OpenStudy (anonymous):

find the exact value of cos(11pi/12)

OpenStudy (lilg132):

-0.96592582628906828674974319972889736763390483900840

OpenStudy (anonymous):

i need the exact answer, not the decimal equivalent

OpenStudy (lilg132):

\[-\frac{ 1+\sqrt{3} }{ 2\sqrt{2} }\]

OpenStudy (anonymous):

cos(11pi/12)=cos(pi -pi/12)=-cos(pi)/12=-cos(pi/4 -pi/3) use cos(a-b)=cosacosb +sinasinb thus cos(11pi/12)=-(sqrt(3)+1)/(2*sqrt(2)) now using calculator u can get ur desired results